OpenStudy (blackops2luvr):

The legs of a right triangle are 3 units and 6 units. What is the length of the hypotenuse? Round your answer to the nearest hundredth. 5.20 units 6.71 units 9.00 units 18.00 units

OpenStudy (blackops2luvr):

@camerondoherty

OpenStudy (mathmate):

Hint: Pythagoras Theorem: \(a^2+b^2=c^2\), or \(c=\sqrt{a^2+b^2}\) where c is the hypotenuse.

OpenStudy (blackops2luvr):

All i need is the explanation and the answer...

OpenStudy (mathmate):

The question has supplied the lengths of a and b (the legs of the triangle. You will need to use one of the formulas to find c, the hypotenuse. Post your answer for a check if you wish.